# Bilevel Reinforcement Learning via the Development of Hyper-gradient without Lower-Level Convexity

### v1

@@ Abstract
Bilevel reinforcement learning (RL), which features intertwined two-level problems, has attracted growing interest recently. The inherent non-convexity of the lower-level RL problem is, however, to be an impediment to developing bilevel optimization methods. By employing the fixed point equation associated with the regularized RL, we characterize the hyper-gradient via fully first-order information, thus circumventing the assumption of lower-level convexity. This, remarkably, distinguishes our development of hyper-gradient from the general AID-based bilevel frameworks since we take advantage of the specific structure of RL problems. Moreover, we propose both model-based and model-free bilevel reinforcement learning algorithms, facilitated by access to the fully first-order hyper-gradient. Both algorithms are provable to enjoy the convergence rate $\mathcal{O}\left({\epsilon^{-1}}\right)$. To the best of our knowledge, this is the first time that AID-based bilevel RL gets rid of additional assumptions on the lower-level problem. In addition, numerical experiments demonstrate that the hyper-gradient indeed serves as an integration of exploitation and exploration.
@@ 1 Introduction
Bilevel optimization, aiming to solve problems with a hierarchical structure, achieves success in a wide range of machine learning applications, e.g., hyper-parameter optimization [19, 20, 83], meta-learning [7], computer vision [49], neural architecture search [47, 78], adversarial training [77], reinforcement learning [32, 10, 69], data poisoning [54]. To address the bilevel optimization problems, a line of research has emerged recently [23, 48, 50, 38, 33, 40, 68, 31, 82]. Generally, the upper-level (resp. the lower-level) problem optimizes the decision taken by the leader (resp. the follower), which exhibits potential for handling complicated decision-making processes such as Markov decision processes (MDPs).
Reinforcement learning (RL) [75, 76] serves as an effective way of learning to make sequential decisions in MDPs, and has seen plenty of applications [70, 6, 59, 58, 74]. The central task of RL is to find the optimal policy that maximizes the expected cumulative rewards in an MDP. Bilevel RL enriches the framework of RL by considering a two-level problem: the follower solves a standard RL problem within an environment parameterized by the decision variable taken by the leader; meanwhile, the leader optimizes the decision variable based on the response policy from the lower level. Recently, bilevel RL has gained increasing attention in practice, including RL from human feedback, [14], inverse RL [8], and reward shaping [34].
@@
where $\Pi$ is the policy set of interest, the upper-level function $f$ is defined on $\mathbb{R}^{n}\times\Pi$, the univariate function $\phi$ is called the hyper-objective, and the gradient of $\phi(x)$ is referred to as the hyper-gradient [63, 24, 11, 80] if it exists. The approximate implicit differentiation (AID) based method which resorts to the hyper-gradient has become flourishing recently [23, 37, 3, 16, 32, 52, 21]. Specifically, in each outer iteration of the AID-based method, one implements an inexact hyper-gradient descent step $x_{k+1}=x_{k}-\beta\widetilde{\nabla}\phi(x_{k})$, where the estimator ~$\widetilde{\nabla}\phi(x_{k})$ of the hyper-gradient is obtained with the help of a few inner iterations. We consider the extension of AID-based methods to the bilevel reinforcement learning problem (1). Note that AID-based methods depend on the lower-level strong convexity [39, 36, 45] or uniform Polyak–Łojasiewicz (PL) condition [35, 10] to ensure the existence of the hyper-gradient. However, the lower-level problem in (1)—always an RL problem—is inherently non-convex even with strongly-convex regularization [1, 41], and only the non-uniform PL condition has been established [57], which renders $\nabla\phi$ ambiguous, as stated in [69].
@@
Recently, Chakraborty et al. [10] presented an AID-based bilevel RL framework, assuming that the lower-level problem satisfies the uniform PL condition and the Hessian non-singularity condition. Shen et al. [69] proposed a penalty-based bilevel RL method to bypass the requirement of lower-level convexity by constructing two penalty functions. The convergence rate relies on the penalty parameter $\lambda$ that is at least the order of ${\epsilon^{-0.5}}$. In this paper, we develop both model-based and model-free bilevel RL algorithms, which are provable and exhibit an enhanced convergence rate, without additional assumptions on the lower-lower problem; see Table 1 for a detailed comparison with the existing bilevel RL methods.
Algorithm Lower-level Assumpsion Conv. Rate Inner Iter. Oracle
PARL [10] PL + Non-singular Hessian $\mathcal{O}(\epsilon^{-1})$      - 1st+2nd
PBRL [69] No $\mathcal{O}(\lambda\epsilon^{-1})$ $\mathcal{O}\left({\log{\lambda^{2}\epsilon^{-1}}}\right)$ 1st
M-SoBiRL (this work) No $\mathcal{O}(\epsilon^{-1})$ $\mathcal{O}\left({1}\right)$ 1st
SoBiRL (this work) No $\mathcal{O}(\epsilon^{-1})$ $\mathcal{O}\left({\log{\epsilon^{-1}}}\right)$ 1st
Contributions.
The main contributions are summarized as follows.
@@
Firstly, we characterize the hyper-gradient of bilevel RL problem via fully first-order information and unveil its properties by investigating the fixed point equation associated with the entropy-regularized RL problem [62, 22, 81], which extends the spirits in [15, 24, 25, 26].
@@
Secondly, understanding the hyper-gradient enables us to construct its estimators, upon which we devise a model-based bilevel RL algorithm, M-SoBiRL, together with a model-free version, SoBiRL. Specifically, the implementation only requires first-order oracles, which circumvents complicated second-order queries in general AID-based bilevel methods. To the best of our knowledge, it is the first time that AID-based bilevel RL algorithms are proposed without additional assumptions on the lower-level problem.
@@
Finally, we offer an analysis to illustrate the efficiency of amortizing the hyper-gradient approximation through outer iterations in M-SoBiRL, i.e., it enjoys the convergence rate $\mathcal{O}(\epsilon^{-1})$ with the inner iteration number $N=\mathcal{O}(1)$ independent of the solution accuracy $\epsilon$. In the model-free scenario, we also establish an enhanced convergence property; refer to Table 1 for detailed results. In addition, the favorable performance of the proposed SoBiRL is validated on the Atari game, which implies that the hyper-gradient is an aggregation of exploitation and exploration.
2 Related Works
@@
Entropy-regularized reinforcement learning. The entropy regularization is commonly considered in RL community. Specifically, the goal of entropy-regularized RL is to maximize the expected reward augmented with the policy entropy, thereby boosting both task success and behavior stochasticity. It facilitates exploration and robustness [87, 28, 29], smoothens the optimization landscape [2], and enhances the convergence property [61, 9, 84, 41, 80, 43]. Moreover, the policy optimality condition under entropy-regularized setting is equivalent to the softmax temporal value consistency in [62]. Therefore, the term soft is prefixed to quantities in this scenario, also resonating with the concept in [75], where “soft” means that the policy ensures positive probabilities across all state-action pairs.
Bilevel reinforcement learning. Extensive reinforcement learning applications hold the bilevel structure, e.g., reward shaping [72, 86, 34, 17, 27], offline reward correction [44], preference-based RL [14, 42, 64, 69], apprenticeship learning [4], Stackleberg Games [18, 71, 69], etc. In terms of provable bilevel reinforcement learning frameworks, Chakraborty et al. [10] proposed a policy alignment algorithm demonstrating performance improvements, and Shen et al. [69] designed a penalty-based method. Nevertheless, the research related to the hyper-gradient in bilevel RL remains constrained.
@@ 3 Problem Formulation
In this section, we will introduce the setting of entropy-regularized Markov decision processes [79, 87, 62, 28], and the formulation of bilevel reinforcement learning. Moreover, two specific instances of bilevel RL are introduced. The notation is listed in Appendix B.
@@ 3.1 Entropy-regularized MDPs
Discounted infinite-horizon MDPs. An MDP is characterized by a tuple $\mathcal{M}_{\tau}=(\mathcal{S},\mathcal{A},P,r,\gamma,\tau)$ with $\mathcal{S}$ and $\mathcal{A}$ serving as the state space and the action space, respectively. In this paper, we restrict the focus to the tabular setting, where both $\mathcal{S}$ and $\mathcal{A}$ are finite, i.e., $\left|{\mathcal{S}}\right|<+\infty$ and $\left|{\mathcal{A}}\right|<+\infty$. Furthermore, $P\in\mathbb{R}^{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|\times\left|{\mathcal{S}}\right|}$ is the transition matrix with $P_{sas^{\prime}}=P(s^{\prime}|s,a)$ representing the transition probability from state $s$ to state $s^{\prime}$ under action $a$. The vector $r\in\mathbb{R}^{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|}$ specifies the reward $r_{sa}$ received when action $a$ is carried out at state $s$. Note that $\gamma\in(0,1)$ is the discount factor, adjusting the importance of immediate versus future reward, and the temperature parameter $\tau\geq 0$ balances regularization and reward.
Soft value function and Q-function. A policy $\pi\in\mathbb{R}^{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|}$ provides an action selection rule, that is, for any $s\in\mathcal{S},\ a\in\mathcal{A}$, $\pi_{sa}=\pi(a|s)$ is the probability of performing action $a$ at state $s$, and we denote $\pi_{s}\in\mathbb{R}^{\left|{\mathcal{A}}\right|}$ to be the corresponding distribution over $\mathcal{A}$ at state $s$, which means $\pi_{s}$ belongs to the probability simplex $\Delta(\mathcal{A}):=\{x\in\mathbb{R}^{\left|{\mathcal{A}}\right|}:x_{i}\geq 0,\ \sum_{i=1}^{\left|{\mathcal{A}}\right|}x_{i}=1\}$. Collecting $\pi_{s}$ over $s\in\mathcal{S}$ defines $\Delta^{\left|{\mathcal{S}}\right|}(\mathcal{A}):=\{\pi=(\pi_{s})_{s\in\mathcal{S}}\in\mathbb{R}^{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|}:\pi_{s}\in\Delta(\mathcal{A})\text{ for }s\in\mathcal{S}\}$ as the feasible set of policies. For simplicity, the above two sets are abbreviated as $\Delta$ and $\Delta^{\left|{\mathcal{S}}\right|}$, respectively. A policy $\pi$ induces a $P^{\pi}\in\mathbb{R}^{\left|{\mathcal{S}}\right|\times\left|{\mathcal{S}}\right|}$, measuring the transition probability from one state to another, i.e., $P^{\pi}_{ss^{\prime}}=\sum_{a}\pi_{sa}P_{sas^{\prime}}$. To promote stochasticity and discourage premature convergence to sub-optimal policies [2], it commonly resorts to the entropy function $h(\cdot)$,
@@
where $\mathbb{R}_{+}^{m}=\{z\in\mathbb{R}_{+}^{m}:z_{i}>0\}$. Consequently, the soft value function $V^{\pi}\in\mathbb{R}^{\left|{\mathcal{S}}\right|}$ represents the expected discounted reward augmented with policy entropy, i.e.,
$V^{\pi}_{s}:=\mathbb{E}\left[{\sum_{t=0}^{\infty}\gamma^{t}\left({r_{s_{t}a_{t}}+\tau h(\pi_{s_{t}})}\right)\ \Big{|}\ s_{0}=s,\pi,\mathcal{M}_{\tau}}\right],\ \text{ for }s\in\mathcal{S},$
where the expectation is taken over the trajectory $(s_{0},a_{0}\sim\pi(\cdot|{s_{0}}),s_{1}\sim P(\cdot|s_{0},a_{0}),\ldots)$. Given the initial state distribution $\rho\in\mathbb{R}_{+}^{\left|{\mathcal{S}}\right|}\cap\Delta\left({{\mathcal{S}}}\right)$, the objective of RL is to find the optimal policy $\pi\in\Delta^{\left|{\mathcal{S}}\right|}$ that solves the problem:
$\max_{\pi\in\Delta^{\left|{\mathcal{S}}\right|}}\ \ \ V^{\pi}(\rho):=\mathbb{E}_{s\sim\rho}\left[{V^{\pi}_{s}}\right]=\mathbb{E}\left[{\sum_{t=0}^{\infty}\gamma^{t}\left({r_{s_{t}a_{t}}+\tau h(\pi_{s_{t}})}\right)\ \Big{|}\ s_{0}\sim\rho,\pi,\mathcal{M}_{\tau}}\right].$ (2)
@@
$\forall(s,a)\in\mathcal{S}\times\mathcal{A}:\quad Q^{\pi}_{sa}\!$ $=r_{sa}+\gamma\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}\left[V^{\pi}_{s^{\prime}}\right],$ (3)
$\forall s\in\mathcal{S}:\quad V^{\pi}_{s}$ $=\mathbb{E}_{a\sim\pi(\cdot|s)}\left[-\tau\log\pi(a|s)+Q^{\pi}_{sa}\right].$
It is worth noting that the optimization problem (2) admits a unique optimal soft policy $\pi^{*}$ independent of $\rho$ [62]. The corresponding optimal soft value function (resp. optimal soft Q-function) is denoted by $V^{*}:=V^{\pi^{*}}$ (resp. $Q^{*}:=Q^{\pi^{*}}$). If necessary, we add a subscript to clarify the environment in which the policy is evaluated, e.g., $V^{\pi}_{\mathcal{M}_{\tau}}$ and $Q^{\pi}_{\mathcal{M}_{\tau}}$.
3.2 Bilevel reinforcement learning formulation
The single-level reinforcement learning task in  Section 3.1 trains an agent with fixed rewards. In contrast, this work considers the scenario where the reward function is parameterized by a decision variable, a task dubbed “policy alignment” in [10]. Generally, we parameterize the reward $r(x)\in\mathbb{R}^{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|}$ with $x\in\mathbb{R}^{n}$, which results in a parameterized MDP $\mathcal{M}_{\tau}(x)=(\mathcal{S},\mathcal{A},P,r(x),\gamma,\tau)$. In parallel with Section 3.1, given $\mathcal{M}_{\tau}(x)$, we define the soft value function $V^{\pi}(x)$, the soft Q-function $Q^{\pi}(x)$, the objective $V_{\mathcal{M}_{\tau}\left(x\right)}^{\pi}\left(\rho\right)$, the optimal soft policy $\pi^{*}(x)$, the optimal soft value-function $V^{*}(x)$, and the optimal soft Q-function $Q^{*}(x)$, associated with $x$.

### v2

@@ Abstract
Bilevel reinforcement learning (RL), which features intertwined two-level problems, has attracted growing interest recently. The inherent non-convexity of the lower-level RL problem is, however, to be an impediment to developing bilevel optimization methods. By employing the fixed point equation associated with the regularized RL, we characterize the hyper-gradient via fully first-order information, thus circumventing the assumption of lower-level convexity. This, remarkably, distinguishes our development of hyper-gradient from the general AID-based bilevel frameworks since we take advantage of the specific structure of RL problems. Moreover, we design both model-based and model-free bilevel reinforcement learning algorithms, facilitated by access to the fully first-order hyper-gradient. Both algorithms enjoy the convergence rate $\mathcal{O}\left({\epsilon^{-1}}\right)$. To extend the applicability, a stochastic version of the model-free algorithm is proposed, along with results on its iteration and sample complexity. In addition, numerical experiments demonstrate that the hyper-gradient indeed serves as an integration of exploitation and exploration.
@@ 1 INTRODUCTION
Bilevel optimization, aiming to solve problems with a hierarchical structure, achieves success in a wide range of machine learning applications, e.g., hyper-parameter optimization (Franceschi et al.,, 2017, 2018; Ye et al.,, 2023), meta-learning (Bertinetto et al.,, 2019), computer vision (Liu et al., 2021a, ), neural architecture search (Liu et al.,, 2019; Wang et al.,, 2022), adversarial training (Wang et al.,, 2021), reinforcement learning (Chakraborty et al.,, 2024; Shen et al.,, 2024; Thoma et al.,, 2024), data poisoning (Liu et al., 2024b, ). To address the bilevel optimization problems, a line of research has emerged recently (Ghadimi and Wang,, 2018; Liu et al.,, 2020; Liu et al., 2021b, ; Ji et al.,, 2022; Hu et al.,, 2023; Kwon et al.,, 2023; Shen and Chen,, 2023; Hao et al.,, 2024; Yao et al.,, 2024). Generally, the upper-level (resp. the lower-level) problem optimizes the decision taken by the leader (resp. the follower), which exhibits potential for handling complicated decision-making processes such as Markov decision processes (MDPs).
Reinforcement learning (RL) (Sutton and Barto,, 2018; Szepesvári,, 2022) serves as an effective way of learning to make sequential decisions in MDPs, and has seen plenty of applications (Silver et al.,, 2017; Berner et al.,, 2019; Mirhoseini et al.,, 2021; Miki et al.,, 2022; Sun et al.,, 2024). The central task of RL is to find the optimal policy that maximizes the expected cumulative rewards in an MDP. Bilevel RL enriches the framework of RL by considering a two-level problem: the follower solves a standard RL problem within an environment parameterized by the decision variable taken by the leader; meanwhile, the leader optimizes the decision variable based on the response policy from the lower level. Recently, bilevel RL has gained increasing attention in practice, including RL from human feedback, (Christiano et al.,, 2017), inverse RL (Brown et al.,, 2019), and reward shaping (Hu et al.,, 2020).
Algorithm Deter. or Stoch. Conv. Rate Inner Iter. Oracle
PARL (Chakraborty et al.,, 2024) Deter. $\mathcal{O}(\epsilon^{-1})$ $\mathcal{O}\left({\log{\epsilon^{-1}}}\right)$ 1st$+$2nd
PBRL (Shen et al.,, 2024) Deter. $\mathcal{O}(\lambda\epsilon^{-1})$ $\mathcal{O}\left({\log{\lambda^{2}\epsilon^{-1}}}\right)$ 1st
HPGD (Thoma et al.,, 2024) Stoch. $\mathcal{O}(\epsilon^{-2})$ $\mathcal{O}\left({\log{\epsilon^{-1}}}\right)$ 1st
M-SoBiRL (this work) Deter. $\mathcal{O}(\epsilon^{-1})$ $\mathcal{O}\left({1}\right)$ 1st
SoBiRL (this work) Deter. $\mathcal{O}(\epsilon^{-1})$ $\mathcal{O}\left({\log{\epsilon^{-1}}}\right)$ 1st
Stoc-SoBiRL (this work) Stoch. $\widetilde{\mathcal{O}}(\epsilon^{-1.5})$ $\mathcal{O}\left({\log{\epsilon^{-1}}}\right)$ 1st
@@
where $\Pi$ is the policy set of interest, the upper-level function $f$ is defined on $\mathbb{R}^{n}\times\Pi$, the univariate function $\phi$ is called the hyper-objective, and the gradient of $\phi(x)$ is referred to as the hyper-gradient (Pedregosa,, 2016; Grazzi et al.,, 2020; Yang et al.,, 2023; Chen et al.,, 2024) if it exists. The approximate implicit differentiation (AID) based method which resorts to the hyper-gradient has become flourishing recently (Ghadimi and Wang,, 2018; Ji et al.,, 2021; Arbel and Mairal,, 2022; Dagréou et al.,, 2022; Liu et al.,, 2023; Hu et al.,, 2024; Yang et al.,, 2025). Specifically, in each outer iteration of the AID-based method, one implements an inexact hyper-gradient descent step $x_{k+1}=x_{k}-\beta\widetilde{\nabla}\phi(x_{k})$, where the estimator ~$\widetilde{\nabla}\phi(x_{k})$ of the hyper-gradient is obtained with the help of a few inner iterations. We consider the extension of AID-based methods to the bilevel reinforcement learning problem (1). Note that AID-based methods depend on the lower-level strong convexity (Khanduri et al.,, 2021; Huang et al.,, 2022; Li et al.,, 2022) or uniform Polyak–Łojasiewicz (PL) condition (Huang,, 2023; Chakraborty et al.,, 2024) to ensure the existence of the hyper-gradient. However, the lower-level problem in (1)—always an RL problem—is inherently non-convex even with strongly-convex regularization (Agarwal et al.,, 2020; Lan,, 2023), and only the non-uniform PL condition has been established (Mei et al.,, 2020), which renders $\nabla\phi$ ambiguous, as stated in (Shen et al.,, 2024).
@@
Recently, Chen et al., 2022a assumed the convexity of the hyper-objective and proposed a deterministic algorithm for the bilevel RL problem. Chakraborty et al., (2024) presented an AID-based framework, assuming that the lower-level problem satisfies the uniform PL condition and the Hessian non-singularity condition. Shen et al., (2024) proposed a penalty-based bilevel RL algorithm to bypass the requirement of lower-level convexity by constructing two penalty functions. The convergence rate relies on the penalty parameter $\lambda$ that is at least the order of ${\epsilon^{-0.5}}$. Thoma et al., (2024) designed a stochastic bilevel RL method, achieving the convergence rate $\mathcal{O}(\epsilon^{-2})$. In this paper, we develop model-based and model-free bilevel RL algorithms and extend the model-free algorithm to stochastic settings, all of which are provable and exhibit an enhanced convergence rate, without the lower-lower convex assumption; see Table 1 for a detailed comparison with the existing works, where the notation $\widetilde{\mathcal{O}}(\cdot)$ hides logarithmic terms of $\epsilon^{-1}$.
Contributions The main contributions are summarized as follows.
@@
Firstly, we characterize the hyper-gradient of bilevel RL problem via fully first-order information and unveil its properties by investigating the fixed point equation associated with the entropy-regularized RL problem (Nachum et al.,, 2017; Geist et al.,, 2019; Yang et al.,, 2019), which extends the spirits in (Christianson,, 1994; Grazzi et al.,, 2020, 2021, 2023).
@@
Secondly, understanding the hyper-gradient enables us to construct its estimators, upon which we devise a model-based bilevel RL algorithm, M-SoBiRL, together with a model-free version, SoBiRL. For broader applicability, we extend SoBiRL in stochastic settings by designing a sampling scheme to estimate the hyper-gradient and a framework Stoc-SoBiRL aided with a momentum technique to accommodate the stochastic hyper-gradient estimator. Specifically, the implementation only requires first-order oracles, which circumvents complicated second-order queries in general AID-based bilevel methods.
@@
Finally, we offer an analysis to illustrate the efficiency of amortizing the hyper-gradient approximation through outer iterations in M-SoBiRL, i.e., it enjoys the convergence rate $\mathcal{O}(\epsilon^{-1})$ with the inner iteration number $N=\mathcal{O}(1)$ independent of the solution accuracy $\epsilon$. In the model-free scenario, an enhanced convergence property is also established. Moreover, investigating the statistical properties of Stoc-SoBiRL, we show its iteration complexity $\widetilde{\mathcal{O}}(\epsilon^{-1.5})$ and sample complexity $\widetilde{\mathcal{O}}(\epsilon^{-3.5})$; refer to Table 1 for a detailed comparison. To the best of our knowledge, it is the first sample complexity result established for bilevel RL problems. In addition, a synthetic experiment verifies the convergence of M-SoBiRL, and the favorable performance of the proposed SoBiRL is validated on Atari games and Mujoco simulations, which implies that the hyper-gradient is an aggregation of exploitation and exploration.
2 RELATED WORK
@@
Entropy-regularized Reinforcement Learning Entropy regularization is commonly considered in the RL community. Specifically, the goal of entropy-regularized RL is to maximize the expected reward augmented with the policy entropy, thereby boosting both task success and behavior stochasticity. It facilitates exploration and robustness (Ziebart,, 2010; Haarnoja et al.,, 2017; Haarnoja et al., 2018a, ; Lan,, 2021), smoothens the optimization landscape (Ahmed et al.,, 2019), and enhances the convergence property (Mnih et al.,, 2016; Cen et al.,, 2022; Zhan et al.,, 2023; Lan,, 2023; Yang et al.,, 2023; Li et al., 2024b, ). Moreover, the policy optimality condition under entropy-regularized setting is equivalent to the softmax temporal value consistency in (Nachum et al.,, 2017). Therefore, the term soft is prefixed to quantities in this scenario, also resonating with the concept in (Sutton and Barto,, 2018), where “soft” means that the policy ensures positive probabilities across all state-action pairs.
Bilevel Reinforcement Learning Extensive reinforcement learning applications hold the bilevel structure, e.g., reward shaping (Sorg et al.,, 2010; Zheng et al.,, 2018; Hu et al.,, 2020; Devidze et al.,, 2022; Gupta et al.,, 2023), offline reward correction (Li et al.,, 2023), preference-based RL (Christiano et al.,, 2017; Lee et al.,, 2021; Saha et al.,, 2023; Shen et al.,, 2024), apprenticeship learning (Arora and Doshi,, 2021), Stackleberg games (Fiez et al.,, 2020; Song et al.,, 2023; Shen et al.,, 2024), etc. In terms of provable bilevel reinforcement learning frameworks, Chakraborty et al., (2024) proposed a policy alignment algorithm demonstrating performance improvements, Shen et al., (2024) designed a penalty-based method, and Thoma et al., (2024) investigated the problem in the stochastic setting. However, the research related to the hyper-gradient in bilevel RL remains constrained.
@@ 3 PROBLEM FORMULATION
In this section, we will introduce the setting of entropy-regularized Markov decision processes (Williams and Peng,, 1991; Ziebart,, 2010; Nachum et al.,, 2017; Haarnoja et al.,, 2017), and the formulation of bilevel reinforcement learning. Moreover, two specific instances of bilevel RL are introduced. The notation is listed in Appendix B.
@@ 3.1 Entropy-regularized MDPs
Discounted Infinite-horizon MDPs An MDP is characterized by a tuple $\mathcal{M}_{\tau}=(\mathcal{S},\mathcal{A},P,r,\gamma,\tau)$ with $\mathcal{S}$ and $\mathcal{A}$ serving as the state space and the action space, respectively. In this paper, we restrict the focus to the tabular setting, where both $\mathcal{S}$ and $\mathcal{A}$ are finite, i.e., $\left|{\mathcal{S}}\right|<+\infty$ and $\left|{\mathcal{A}}\right|<+\infty$. Furthermore, $P\in\mathbb{R}^{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|\times\left|{\mathcal{S}}\right|}$ is the transition matrix with $P_{sas^{\prime}}=P(s^{\prime}|s,a)$ representing the transition probability from state $s$ to state $s^{\prime}$ under action $a$. The vector $r\in\mathbb{R}^{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|}$ specifies the reward $r_{sa}$ received when action $a$ is carried out at state $s$. Note that $\gamma\in(0,1)$ is the discount factor, adjusting the importance of immediate versus future reward, and the temperature parameter $\tau\geq 0$ balances regularization and reward.
Soft Value Function and Q-function A policy $\pi\in\mathbb{R}^{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|}$ provides an action selection rule, that is, for any $s\in\mathcal{S},\ a\in\mathcal{A}$, $\pi_{sa}=\pi(a|s)$ is the probability of performing action $a$ at state $s$, and we denote $\pi_{s}\in\mathbb{R}^{\left|{\mathcal{A}}\right|}$ to be the corresponding distribution over $\mathcal{A}$ at state $s$, which means $\pi_{s}$ belongs to the probability simplex $\Delta(\mathcal{A}):=\{x\in\mathbb{R}^{\left|{\mathcal{A}}\right|}:x_{i}\geq 0,\ \sum_{i=1}^{\left|{\mathcal{A}}\right|}x_{i}=1\}$. Collecting $\pi_{s}$ over $s\in\mathcal{S}$ defines $\Delta^{\left|{\mathcal{S}}\right|}(\mathcal{A}):=\{\pi=(\pi_{s})_{s\in\mathcal{S}}\in\mathbb{R}^{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|}:\pi_{s}\in\Delta(\mathcal{A})\text{ for }s\in\mathcal{S}\}$ as the feasible set of policies. For simplicity, the above two sets are abbreviated as $\Delta$ and $\Delta^{\left|{\mathcal{S}}\right|}$, respectively. A policy $\pi$ induces a $P^{\pi}\in\mathbb{R}^{\left|{\mathcal{S}}\right|\times\left|{\mathcal{S}}\right|}$, measuring the transition probability from one state to another, i.e., $P^{\pi}_{ss^{\prime}}=\sum_{a}\pi_{sa}P_{sas^{\prime}}$. To promote stochasticity and discourage premature convergence to sub-optimal policies (Ahmed et al.,, 2019), it commonly resorts to the entropy function $h(\cdot)$,
@@
where $\mathbb{R}_{+}^{m}:=\{z\in\mathbb{R}^{m}:z_{i}>0\}$. Consequently, the soft value function $V^{\pi}\in\mathbb{R}^{\left|{\mathcal{S}}\right|}$ represents the expected discounted reward augmented with policy entropy, i.e.,
$V^{\pi}_{s}:=\mathbb{E}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}\left({r_{s_{t}a_{t}}+\tau h(\pi_{s_{t}})}\right)\ \big{|}\ s_{0}=s,\pi\Big{]},$
for $s\in\mathcal{S}$, where the expectation is taken over the trajectory $(s_{0},a_{0}\sim\pi(\cdot|{s_{0}}),s_{1}\sim P(\cdot|s_{0},a_{0}),\ldots)$. Given the initial state distribution $\rho\in\mathbb{R}_{+}^{\left|{\mathcal{S}}\right|}\cap\Delta\left({{\mathcal{S}}}\right)$, the objective of RL is to find the optimal policy $\pi\in\Delta^{\left|{\mathcal{S}}\right|}$ that solves the problem:
$\max_{\pi\in\Delta^{\left|{\mathcal{S}}\right|}}\ \ \ V^{\pi}(\rho):=\mathbb{E}_{s\sim\rho}\left[{V^{\pi}_{s}}\right].$ (2)
@@
$\forall(s,a)\in\mathcal{S}\times\mathcal{A}:Q^{\pi}_{sa}\!$ $=r_{sa}+\gamma\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}\left[V^{\pi}_{s^{\prime}}\right],$ (3)
$\forall s\in\mathcal{S}:V^{\pi}_{s}$ $=\mathbb{E}_{a\sim\pi(\cdot|s)}\left[-\tau\log\pi(a|s)+Q^{\pi}_{sa}\right].$
It is worth noting that the optimization problem (2) admits a unique optimal soft policy $\pi^{*}$ independent of $\rho$ (Nachum et al.,, 2017). The corresponding optimal soft value function (resp. optimal soft Q-function) is denoted by $V^{*}:=V^{\pi^{*}}$ (resp. $Q^{*}:=Q^{\pi^{*}}$). If necessary, we add a subscript to clarify the environment in which the policy is evaluated, e.g., $V^{\pi}_{\mathcal{M}_{\tau}}$ and $Q^{\pi}_{\mathcal{M}_{\tau}}$.
3.2 Bilevel RL Formulation
The single-level reinforcement learning task in  Section 3.1 trains an agent with fixed rewards. In contrast, this work considers the scenario where the reward function is parameterized by a decision variable, a task dubbed “policy alignment” in (Chakraborty et al.,, 2024). Generally, we parameterize the reward $r(x)\in\mathbb{R}^{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|}$ with $x\in\mathbb{R}^{n}$, which results in a parameterized MDP $\mathcal{M}_{\tau}(x)=(\mathcal{S},\mathcal{A},P,r(x),\gamma,\tau)$. In parallel with Section 3.1, given $\mathcal{M}_{\tau}(x)$, we define the soft value function $V^{\pi}(x)$, the soft Q-function $Q^{\pi}(x)$, the objective $V_{\mathcal{M}_{\tau}\left(x\right)}^{\pi}\left(\rho\right)$, the optimal soft policy $\pi^{*}(x)$, the optimal soft value-function $V^{*}(x)$, and the optimal soft Q-function $Q^{*}(x)$, associated with $x$.